### v1

# EasyText: Controllable Diffusion Transformer for Multilingual Text Rendering

Abstract
Generating accurate multilingual text with diffusion models has long been desired but remains challenging. Recent methods have made progress in rendering text in a single language, but rendering arbitrary languages is still an unexplored area. This paper introduces EasyText, a text rendering framework based on DiT (Diffusion Transformer), which connects denoising latents with multilingual character tokens encoded as character tokens. We propose character positioning encoding and position encoding interpolation techniques to achieve controllable and precise text rendering. Additionally, we construct a large-scale synthetic text image dataset with 1 million multilingual image-text annotations as well as a high-quality dataset of 20K annotated images, which are used for pretraining and fine-tuning respectively. Extensive experiments and evaluations demonstrate the effectiveness and advancement of our approach in multilingual text rendering, visual quality, and layout-aware text integration. Project page: https://github.com/songyiren725/EasyText
Figure 1: Text-rendered results generated by EasyText, which supports text rendering in over ten languages and produces high-quality results. It can render text either with explicit positional control or in a layout-free manner, and effectively handles curved and slanted regions. The displayed texts do not appear in the prompts; they are included solely to illustrate the intended rendering targets.
1 Introduction
Scene text rendering plays a crucial role in various real-world applications. However, most existing methods such as TextDiffuser [7, 8], Diff-font [14] and modern commercial models like FlUX-dev [20] and Ideogram [1], are primarily limited to English, making multilingual text rendering still a challenging task. GlyphBy-T5-V2 [25] was the first to enable multilingual text rendering by introducing a specially designed glyph encoder and employing a multi-stage training strategy, achieving text rendering in 10 languages on a large-scale dataset of millions of images.
Inspired by how humans learn to write, we gain several key insights: (1) Imitative writing is significantly easier than recalling—humans typically begin by mimicking (i.e., writing with visual reference) before progressing to memory-based writing. (2) Once familiar with one language, humans naturally develop the ability to reproduce text in other unfamiliar languages even without understanding them—treating it more like drawing than writing. Motivated by this, we argue that training AI to “imitate” rather than “recall” is a more efficient and effective strategy for text rendering.
The task of scene text rendering faces several key challenges: (1) Multilingual character modeling is highly complex—for example, Chinese alone has over 30,000 characters. Extending to multilingual settings drastically expands the character space, making joint modeling harder, especially for rare or low-frequency characters with poor rendering accuracy. This is further complicated by language imbalance and font variability. (2) Text-background integration is often unnatural; existing methods struggle to blend rendered text with scene content, leading to visual artifacts such as disconnection or pasted-on effects that undermine image realism. (3) Preserving generative priors is difficult, as fine-tuning on large-scale text-image datasets improves rendering capabilities but often degrades the model’s general image generation ability.
Table 1: Functionality evaluation of EasyText in comparison to other competitors
Functionality Position Control Irregular Regions Multi-Lingual Long Text Rendering Text-Image Blending Unseen Characters Generalization
Glyph-SDXL-v2
AnyText
SD3.5
FLUX-dev
Jimeng AI 2.1
EasyText
To this end, we present EasyText (shown in Fig. 1), a novel multilingual text image generation framework based on Diffusion Transformers [30, 46]. We encode text into font tokens via a VAE and concatenate them in the latent space with denoised latents. By leveraging the powerful in-context capabilities of DiT, EasyText achieves high-quality and accurate text rendering. Additionally, we propose a simple yet effective position control strategy called Implicit Character Position Alignment, which allows for precise control of character positions through positional encoding interpolation and replacement—enabling both position-aware rendering and layout-free generation.
EasyText is also highly efficient in terms of data usage. Unlike GlyphBy-T5 V1/V2 [24, 25], which rely on contrastive synthetic data or massive real-world scene text datasets, our method simply overlays text randomly on natural images during the pretraining stage to learn glyph features. In this stage, to enable the model to learn glyph imitating rather than simple shape copying, we employ a multi-font mapping approach. Multiple different fonts are overlaid in the synthetic training images, while the condition image uses only a standard font. Afterwards, we fine-tune a lightweight LoRA [16] on only 20K high-quality multilingual scene text images, enhancing the visual consistency between text and background. As shown in Table 1, we highlight some of the capabilities demonstrated by EasyText including: (a) Position control, precisely positioning text in specific locations within an image; (b) Irregular regions, enabling it to handle text rendering in irregular regions such as slanted or curved regions; (c) Multilingual text handling, supporting text generation in multiple languages such as Chinese, English, Japanese, etc; (d) Long text rendering, which can render extended text passages; (e) Text-image blending, which integrates text seamlessly into images, maintaining natural visual consistency; and (f) Unseen character generalization, allowing the model to exhibit generalization capability on unfamiliar and unseen characters.
In summary, our contributions are:
1.We propose EasyText, a framework that teaches AI to "imitate" rather than "recall," achieving high-quality multilingual text rendering by harnessing the in-context learning power of Diffusion Transformers.
2.We introduce Implicit Character Position Alignment, which precisely controls text placement via lightweight position encoding operations and supports layout-free generation.
3.Extensive experiments demonstrate the effectiveness and simplicity of our method, showing superior performance on challenging scenarios such as long text, multi-text layouts, irregular regions, and unseen characters.
2 Related Works
2.1 Diffusion Models
In recent years, diffusion models [23, 35] have achieved remarkable progress and emerged as a widely used approach for image generation, offering strong capabilities in producing high-quality and diverse visual content through iterative denoising. Their strong capacity for modeling complex data distributions has enabled broad applicability, ranging from image synthesis and editing [4, 15] to video generation [3]. Notable examples include Stable Diffusion [33], its enhanced version SDXL [31], and subsequent variants, which have demonstrated the scalability and effectiveness of text-to-image diffusion models for high-quality image synthesis. Recently, the use of transformer-based denoisers—particularly the Diffusion Transformer (DiT) [30], which replaces U-Net with a transformer backbone—has gained significant traction, and has been adopted in many state-of-the-art models such as FlUX-dev [20], Stable Diffusion 3.5 [11], and PixArt [9].
2.2 Condition-guided Diffusion Models
Condition-guided diffusion models aim to incorporate external signals—such as spatial structure or semantic reference—into the generative process, enabling control over layout [22, 58], content [55, 56, 50, 57, 59, 37], and identity [49]. These methods facilitate consistent synthesis, alignment with user intent, and support for flexible customization. Early implementations based on U-Net architectures adopted two main paradigms: attention-based conditioning, which integrates semantic features via auxiliary encoders, and residual-based fusion, which injects spatial features into intermediate layers. ControlNet [54] and T2I-Adapter [28] are representative examples that improve spatial consistency and layout control under this framework. With the adoption of transformer-based diffusion models, recent approaches reformulate both semantic and spatial conditions as token sequences. These are integrated into the generation process through multi-modal attention or token concatenation mechanisms, as demonstrated in DiT-based systems like OminiControl [41, tan2025ominicontrol2] . This unified formulation advances the development and application of conditional image generation  [47, 38, 36, 17, 13], improving scalability, simplifying model design, and supporting efficient handling of multiple or high-resolution conditions.
2.3 Visual Text Generation
Text generation and rendering is a classic task, where early methods mainly relied on generative adversarial networks (GANs) [51, 6, 2, 42] and vector stroke techniques  [43, 39, 40]. Recent work on visual text generation with diffusion models primarily focuses on optimizing text encoding and spatial control mechanisms to improve the fidelity and controllability of rendered text. Character-aware encoders are increasingly adopted: GlyphDraw [27], GlyphControl [52], and AnyText [45] embed glyph or OCR features into conditional inputs, while TextDiffuser-2 [8] leverages character-level tokenization to improve alignment. However, semantic bias and tokenizer limitations still persist. Spatial control has been addressed by introducing explicit layout-related conditions. TextDiffuser [7] and ControlText [18] use segmentation or layout masks, and UDiffText [60] and Brush Your Text [53] refine attention to enforce region-level alignment. For multilingual rendering, works like AnyText [45], GlyphControl [52], and Glyph-ByT5-v2 [25] bypass tokenizers using glyph encoders or tokenizer-free models. Yet, rendering quality remains limited by encoder–diffusion compatibility. Newer models (GlyphDraw2 [26], JoyType [21], AnyText2 [44], FonTS [34], RepText [48]) improve layout precision, while closed models (Kolors 2.0 [19], Seedream 3.0 [12], GPT-4o [29]) show promising results but often suffer from limited spatial controllability and suboptimal performance in multi-text layouts. Despite progress, challenges remain in complex layouts and multilingual text, motivating further integration of glyph, position, and semantic signals.
Figure 2: Overview of EasyText. We adopt a two-stage training strategy: large-scale pretraining for glyph generation and spatial mapping, followed by fine-tuning for visual-text integration and aesthetic refinement. Character positions from the condition input are aligned with target regions via implicit character position alignment, and training proceeds with image-conditioned LoRA.
3 Method
Section 3.1 introduces FLUX as the base model; Section 3.2 outlines the overall architecture; Section 3.3 describes the target text condition representation; Section 3.4 details the Implicit Character Position Alignment; and Section 3.5 explains the paired dataset construction.
3.1 Preliminary: FLUX
FLUX is an advanced generative framework significantly enhancing generation quality. Departing from traditional U-Net-based diffusion models (such as Stable Diffusion), it adopts a Transformer backbone (DiT) to improve long-sequence modeling and complex semantic understanding. The architecture employs a pre-trained T5 [32] encoder to project textual inputs into dense semantic embeddings $\mathbf{C}_{T}\in\mathbb{R}^{M\times d}$. For visual processing, a variational autoencoder (VAE) performs nonlinear dimensionality reduction on pixel space, generating compact latent representations that preserve essential visual semantics. Additionally, the framework processes encoded noise patches $\mathbf{X}\in\mathbb{R}^{N\times d}$ through rotary position embeddings (RoPE), maintaining spatial coherence during the iterative denoising process. These noise and text tokens are then fused by concatenation to form an image-text token sequence and then fed in parallel into a transformer-based denoising model. The model iteratively denoises and progressively restores a high-quality image aligned with the textual prompt.
3.2 Overall Architecture
The training pipeline adopts a two-stage strategy. The first stage involves training on a large-scale, synthetically generated dataset with balanced multilingual coverage, enabling the model to learn glyph generation across diverse scripts and accurate spatial mapping. In the second stage, we fine-tune the model on a high-quality annotated dataset to improve both the aesthetic quality of rendered text and its integration with complex scene content. For conditioning, the input text image is encoded into the same latent space as the target image via a VAE. Its positional encoding is aligned with the target region through implicit character position alignment, then concatenated with the target image’s position encoding before being passed to the FLUX DiT block. (Shown in Fig. 2) Image-conditioned LoRA adaptation is then applied, with both the VAE and text encoder kept frozen.
3.3 Character Representation in EasyFont
In contrast to conventional font conditioning approaches that typically employ symbolic or parametric representations, our methodology innovatively adopts a visually-grounded paradigm inspired by glyph morphology. The key differentiation of our character representation lies in its image-based conditioning architecture, where discrete image patches serve as fundamental conditioning units. This design philosophy stems from two critical observations:
First, we establish a unified multilingual representation to addresses the intrinsic typographic differences between writing systems. For alphabetic scripts (e.g., English), we use 64-pixel-high images with widths adaptive to text length, naturally capturing the connected structure of alphabetic word formations. This horizontal layout preserves crucial inter-character relationships and spacing conventions unique to Western typography. Second, for logographic systems (Chinese, Japanese, etc.), we allocate individual 64×64 square images per character. This configuration respects the isolated nature of ideographic characters while maintaining consistent resolution across all glyphs. The square aspect ratio optimally captures the balanced structure inherent to CJK characters, where stroke complexity is uniformly distributed within a square design space.
The image-based character representations are encoded via VAE encoder into latent features. Subsequently, the latent representations undergo patchification through a linear projection layer, which reshapes and projects the spatial features into a sequence of conditioning tokens.
3.4 Implicit Character Position Alignment
To achieve precise spatial control in font generation, we propose an Implicit Character Position Alignment (ICPA) mechanism through positional encoding extrapolation. Given a conditional image patch $\mathbf{P}_{c}\in\mathbb{R}^{64\times W_{c}\times 3}$ containing source typography features, $W_{c}$ denotes the total width of the conditional image patch, and a target rendering region $\Omega_{t}=[x_{1}:x_{2},y_{1}:y_{2}]$ in the output image, our method establishes position-aware feature correspondence through positional encoding extrapolation.
Linear Alignment via Affine Transform. Linearly interpolate the position in the condition image (of size $64\times W_{c}$) to the position in the target bounding box $\Omega_{t}=[x_{1}:x_{2},y_{1}:y_{2}]$. Assuming $u\in[0,W_{c}-1]$ and $v\in[0,63]$ (i.e. the condition image spans indices 00 to $W_{c}$ horizontally and 00 to $63$ vertically), the affine mapping $\mathcal{T}_{\text{aff}}:(u,v)\mapsto(x,y)$ is defined as:
$\mathcal{T}_{\text{aff}}(u,v)=\Big{(}x_{1}+\frac{u}{\,W_{c}-1\,}(x_{2}-x_{1}),\quad y_{1}+\frac{v}{\,63\,}(y_{2}-y_{1})\Big{)}\,,$ (1)
which maps the normalized horizontal coordinate $u/(W_{c}-1)$ to a point between $x_{1}$ and $x_{2}$, and similarly $v/63$ to a point between $y_{1}$ and $y_{2}$, achieving a linear scaling and translation of the patch coordinates into the target domain. The positional encoding is then re-aligned by applying the affine transform to each coordinate in condition image $(u,v)$. Specifically, the transformed positional encoding at the target location $\text{PE}_{t}(x,y)=\mathcal{T}_{\text{aff}}(u,v)$ is assigned from the location in condition image. Thus, the new positional encoding for points in the condition image is given by:
$\text{PE}_{c}(u,v)=\text{PE}_{t}(x,y),$ (2)
where the target position $(x,y)$ is computed via the affine mapping from the patch coordinates.
Nonlinear Alignment via Thin-Plate Spline Interpolation. For nonlinear spatial alignment, we use a Thin-Plate Spline (TPS) interpolation. TPS warping defines a smooth map that exactly interpolates a set of control point correspondences between the source patch and target region, while minimizing bending (second-derivative) distortion. Let $\{(u_{i},v_{i})\}_{i=1}^{K}$ be $K$ landmark points in the source patch $P_{c}$ and $\{(x_{i},y_{i})\}_{i=1}^{K}$ their corresponding positions in the target region $\Omega_{t}$. The TPS mapping $\mathcal{T}_{\text{TPS}}:(u,v)\mapsto(x,y)$ can be expressed as an affine base plus a radial basis deformation term:
$\mathcal{T}_{\text{TPS}}(u,v)=A\begin{pmatrix}u\\ v\\ 1\end{pmatrix}+\sum_{i=1}^{K}w_{i}\,\phi\!\left(\sqrt{(u-u_{i})^{2}+(v-v_{i})^{2}}\right),\quad\phi(r)=r^{2}\log r,$ (3)
where $A\in\mathbb{R}^{2\times 3}$ represents the affine part of the transformation and $w_{i}\in\mathbb{R}^{2}$ are the TPS warp coefficients associated with each control point. The kernel $\phi(r)=r^{2}\ln r$ is the fundamental solution of the biharmonic equation in 2D. The parameters $A$ and ${w_{i}}$ are determined by solving the interpolation constraints
$T_{\text{TPS}}(u_{i},v_{i})=(x_{i},y_{i}),i=1,\dots,K,$ (4)
together with additional conditions to ensure a well-posed solution ($\sum_{i=1}^{K}w_{i}=0$, $\sum_{i=1}^{K}w_{i}u_{i}=0$, $\sum_{i=1}^{K}w_{i}v_{i}=0$). This yields a smooth, non-linear warp $\mathcal{T}_{\text{TPS}}$ that maps every source coordinate $(u,v)$ to a target coordinate $(x,y)$, honoring the given correspondences. Similar to linear alignment, target location is obtained via $\text{PE}_{t}(x,y)=\mathcal{T}_{\text{TPS}}(u,v)$. Accordingly, the corresponding positional encoding $\text{PE}_{c}(u,v)$ can be computed as in Equation 2.
Layout-Free Position Alignment via Positional Offset Injection. To enable flexible layout-free rendering, we also introduce an effective positional offset strategy. Instead of aligning coordinates, we shift the positional encoding of the conditional patch by a fixed scalar offset $w_{t}$, which represents the width of the target image. Let $(i,j)$ index the spatial position in $\mathbf{P}_{c}$. Target coordinates $(x^{\prime},y^{\prime})$ are obtained by applying a positional offset to the original coordinates:
$\begin{cases}x^{\prime}=i+w_{t}\\ y^{\prime}=j\end{cases}$ (5)
This ensures positional encoding uniqueness without binding the conditional patch to any specific target location, enabling more flexible text rendering.
3.5 EasyText Dataset Construction
We construct two datasets tailored to the specific needs of the pretraining and fine-tuning stages.
Large-Scale Synthetic Dataset. This dataset is designed to provide broad coverage across scripts and languages. Target images are generated by rendering multilingual text onto background images using script-based synthesis. The dataset includes scripts such as Latin, Chinese, Korean, and others, resulting in nearly 1 million samples. To reduce redundancy and improve efficiency, languages sharing the same writing system are grouped by script—for example, English, Italian, and German are treated as a single Latin script class. The rendered text is composed of random combinations of characters from the corresponding script, covering its full character set. To encourage the model to learn generalizable glyph representations rather than simply copying shapes from the condition image, the target text is rendered using diverse fonts, while the condition image is rendered using a standardized font. This font decoupling enables robust learning of character structure across styles. Dataset statistics are summarized in Table 3.
High-Quality Human-Annotated Dataset. The second dataset comprises 20k high-quality image–text box pairs, primarily in Chinese and English. Text regions are extracted using PP-OCR [10] and filtered to ensure annotation quality. In these examples, text is more naturally integrated into scene content, often with stylized typography and visually consistent stroke patterns. The rendered text is not included explicitly in the prompt, but instead triggered by placeholder tokens (e.g., <sks1>, <sks2>), supporting semantic alignment between text and background. This dataset facilitates better visual-text fusion and improves typographic aesthetics.
This two-stage training strategy reduces the need for large-scale text-image datasets, requiring only a small fine-tuning set—particularly well-suited in multilingual settings where annotated data is scarce.
4 Experiments
4.1 Experimental Settings
Implementation Details. Our method is implemented based on the open-source FLUX-dev framework with LoRA-based parameter-efficient tuning. In the first stage, we set the LoRA rank to 128 and train on 4 H20 GPUs. The model is first trained at 512×512 resolution for 48,000 steps with a batch size of 24, and then at 1024×1024 for 12,000 steps. In the second stage, we reduce the LoRA rank to 32 and fine-tune at 1024 resolution for 10,000 steps with a batch size of 8. We use the AdamW optimizer with a learning rate of 1e-4.
Evaluation Metric. To evaluate the accuracy of multilingual text generation, we adopt two complementary metrics: (i) Character-level precision which measures the correctness of generated characters relative to ground-truth texts. This reflects the model’s ability to produce accurate predictions at the most basic unit level for each language. (ii) Sentence-level precision, which assesses whether the entire text line content within each controlled text box is rendered completely correctly. This captures the overall consistency and correctness at the region level.
Multilingual Benchmark. We construct a multilingual benchmark with 90 language-agnostic prompts across 10 languages. For each language, prompts are paired with language-specific text to generate condition images while preserving semantic intent. Four images are generated per prompt-language pair. For logographic scripts (Chinese, Japanese, Korean), the average number of characters per text box is 7.1, and for alphabetic scripts, it is 14.3. Each image contains an average of 1.7 text boxes. To evaluate the generated results, three volunteers were tasked with assessing the accuracy of the text in the images, ensuring reliable multilingual generation quality.
Table 2: Statistics of multilingual data used in large-scale pre-training and fine-tuning.
Script Type Large-Scale Pre-Training Fine-Tuning
Unique Chars Font types Samples Samples
Chinese 7000 18 230K 5.5K
Korean 4308 13 100K 0.1K
Japanese 2922 17 100K 0.1K
Thai 2128 3 100K
Vietnamese 91 19 100K
Latin 104 30 180K 16K
Greek 79 19 100K
Table 3: Character/Sentence level precision for English and Chinese across different methods.
Method English Chinese
Chars Pre Sen Pre Chars Pre Sen Pre
EasyText(pretrain) 0.9968 0.9813 0.9344 0.6562
EasyText(fine-tune) 0.9945 0.9625 0.9312 0.6438
Glyph-SDXL-v2 0.9959 0.9688 0.9258 0.6250
FLUX 0.9180 0.7062
Stable Diffusion 3.5 0.9556 0.7938
Jimeng 2.1 0.9812 0.8687 0.9214 0.6813
ANYTEXT 0.8978 0.6364 0.8824 0.6071
Table 4: Multilingual text generation precision with our unified and generalized model, compared to Glyph-SDXL-v2.
Language EasyText (pre) EasyText (fine) Glyph-ByT5-v2
Chars Sent Chars Sent Chars Sent
French 0.9867 0.8552 0.9673 0.7500 0.9812 0.8625
German 0.9818 0.8333 0.9732 0.7619 0.9828 0.7976
Korean 0.9341 0.7125 0.9203 0.6713 0.9441 0.7437
Japanese 0.9265 0.7179 0.9194 0.6562 0.9059 0.6187
Italian 0.9740 0.9125 0.9638 0.8571 0.9385 0.8333
Thai 0.9628 0.7443 0.9334 0.6500
Vietnamese 0.9605 0.7312 0.9401 0.6474
Greek 0.9702 0.7685 0.9360 0.6875
Table 5: Quantitative evaluation of text-to-image rendering across diverse metrics and GPT-4o-based assessments.Ours (pre) and Ours (ft) denote the pretrained and finetuned versions of our model, respectively.
Metric Ours(pre) Ours(ft) Glyph-v2 SD3.5 FLUX
CLIPScore 0.3318 0.3486 0.3140 0.3519 0.3348
OCR Acc.(%) 84.32 88.72 82.33 79.33 76.09
Evaluation by GPT-4o
Image Aesthetics 81.58 83.86 75.86 84.58 83.61
Text Aesthetics 65.14 73.79 65.28 72.06 71.90
Text Quality 84.58 90.66 86.25 85.79 86.95
Text-Image Fusing 74.48 80.28 74.80 81.12 78.66
Figure 3: Qualitative comparison of EasyText with other methods, focusing on the generation quality of both text and images, reveals that EasyText demonstrates outstanding performance.
Figure 4: Qualitative comparison of EasyText across multiple languages of the same prompt.
4.2 Comparision results
Multi-lingual Precision Evaluation. We evaluate our model on a multilingual benchmark against state-of-the-art commercial models (e.g., FLUX [20], Jimeng AI [5], SD3.5 [11]) and other methods (e.g., Glyph-ByT5-v2 [25], AnyText [45]), following the benchmark and metrics outlined in Sec. 4.1. For languages with limited support in prior rendering methods, we compare primarily with Glyph-SDXL-v2. All baseline models are evaluated using their official inference settings. For each sample, four images are generated with identical image descriptions and target texts across all methods. For models lacking conditional input support, the target text is included in the prompt; Evaluation focuses solely on text accuracy, regardless of spatial layout or surrounding content.
Figure 5: Comparison of text rendering on the pretraining synthetic dataset using multiple font overlays versus single font overlays. The red boxes highlight the erroneous regions.
Comprehensive Quality Assessment. Beyond precision evaluation, we assess generation quality using CLIPScore and OCR Accuracy for objective comparison against existing baselines. We further incorporate GPT-4o assessment and a user study across four subjective criteria: Image Aesthetics, Text Aesthetics, Text Quality, and Text-Image Fusing, to evaluate overall fidelity and alignment.
The results show that our model surpasses competing methods in text rendering precision across several languages, including English and Italian (Table 5) and achieves strong character-level accuracy in Chinese, though slightly behind Jimeng AI at the sentence level. It also performs well in unsupported languages like Thai and Greek. In Table 5, it also leads in OCR accuracy and improves over FLUX in CLIPScore, indicating higher visual-text alignment. These trends are further supported by GPT-4o evaluation results. After pretraining, the model demonstrates strong text rendering performance from condition images. However, it shows limited text-image coherence, as indicated by lower CLIPScore and GPT-4o evaluations. Fine-tuning alleviates this issue, significantly improving CLIPScore, Text Aesthetics, and overall visual-text alignment. Although rendering accuracy slightly declines due to flexible positioning and artistic font usage, the trade-off is acceptable given the qualitative gains. Notably, while most competing methods use datasets exceeding 1M samples, our fine-tuning relies on only  20k annotated examples. Moreover, the large-scale pretraining set can be replaced with synthetic data, reducing dependence on large, annotated datasets.
4.3 Qualitative Results
We provide qualitative comparisons of multilingual text rendering, including Chinese, English, and other languages. Results are shown in Fig. 4. Compared to previous region-controlled text rendering methods, our approach demonstrates significant improvements in visual quality and text fidelity, with better visual-text integration, higher OCR accuracy, and enhanced aesthetic coherence (Shown in Fig. 3). Additionally, unlike commercial text-to-image models, our method allows for more precise spatial control over rendered text, particularly in generating multiple paragraphs of relatively long text (e.g., 4–5 paragraphs with around 20 characters each) while maintaining consistent layouts. We further evaluate the model’s ability to handle challenging text generation scenarios that are typically difficult for existing methods. Specifically, our approach generalizes well to unseen characters, slanted or non-rectilinear text regions, and long text sequences, while maintaining structural consistency and legibility. Detailed qualitative results are provided in the appendix.
Figure 6: Ablation study comparing the full conditioned method with: (1) layout-free (2) without condition inputs, and (3) original FLUX. The red boxes highlight the erroneous regions.
4.4 Ablation Studies
To validate the effectiveness of using diverse fonts in synthetic pretraining, we conducted a controlled experiment where both the target and condition images were rendered using the same standard font. While this setup yields high initial precision, it mainly captures a direct shape mapping and fails to generalize after fine-tuning on real-world data(Shown in Fig. 5). This highlights the importance of font diversity—rendering targets with multiple fonts while keeping the condition font fixed—for learning robust, transferable representations. we compare our full approach—using both the condition input and implicit character position alignment-with three ablated variants: (1) removing position alignment, (2) removing the condition input, and (3) using the original FLUX model. As shown in Fig. 6, incorporating position mapping substantially improves spatial precision and text accuracy, particularly in multi-text scenarios. In comparison, when removing the condition input, the target render text is instead provided in the prompt(in our full method, renter text content is given via the condition input). This demonstrates that our fine-tuned model effectively retains the strong generation capability of the base FLUX model.
5 Limitation
The Implicit Character Position Alignment mechanism is less effective when character positions are substantially overlapping, occasionally leading to reduced rendering accuracy. In addition, training across multiple scripts results in confusions between simple but visually similar characters from different writing systems, such as rendering the Vietnamese character "" as "e". These cases are infrequent but consistently observed.
6 Conclusion
In this work, we propose EasyText, a diffusion-based framework for multilingual and controllable text generation in text-to-image synthesis. The method learns to mimic font rendering by mapping glyph features into the same latent space as the image via a VAE, and achieves Implicit Character Position Alignment by combining position encoding with spatial targets, enabling precise and flexible text placement. Benefiting from a two-stage training strategy—pretraining on a large-scale synthetic multi-font dataset followed by fine-tuning—EasyText reduces the need for real multilingual data while maintaining high rendering accuracy. It also demonstrates strong visual-text integration, effectively embedding text into complex scenes. Extensive experiments confirm the effectiveness and generalization ability of our approach.
Supplementary Material for EasyText: Controllable Diffusion Transformer for Multilingual Text Rendering
David S. HippocampusUse footnote for providing further information about author (webpage, alternative address)—not for acknowledging funding agencies.
Department of Computer Science
Cranberry-Lemon University
Pittsburgh, PA 15213
hippo@cs.cranberry-lemon.edu
Appendix
This supplementary material is organized into three sections. Section A describes the training datasets used for pretraining and fine-tuning, including illustrative examples. Section 12 presents diverse and comprehensive generated examples. Section C outlines the process and sample cases of the GPT-based evaluation and user study.
Appendix A Additional Dataset Details
The main paper provides a detailed description of the dataset’s design and composition. This section complements that discussion by offering additional details and presenting example instances from the dataset to further illustrate its structure.
Figure 7: Sample images from the synthetic dataset used in the pretraining.
A.1 Synthetic Dataset for Large-Scale Pretraining
For the training dataset used in pretraining, we collected background images from several publicly available datasets without textual content, such as LAION-Aesthetics. For each script (i.e., a language-specific character set), we rendered text in various fonts onto these backgrounds. During the rendering process, we recorded both the text content and the corresponding bounding box coordinates, producing synthetic image–text pairs designed for text rendering tasks. This approach enabled the creation of a large-scale multilingual dataset of approximately 1 million samples, offering a practical alternative for low-resource languages, where high-quality annotated image–text pairs are often unavailable. Some sample instances are shown in Fig. 7.
A.2 High-Quality Dataset for Fine-Tuning
For the fine-tuning stage, we collected around 20K images containing text from copyright-free websites. Text regions were detected and filtered using PP-OCR, and each image was further annotated with a caption generated by GPT-4 Mini. Some sample instances are shown in Fig. 8.
Figure 8: Sample images from the high-quality dataset used in the fine-tuning.
Appendix B Additional Generation Results
We present additional diverse generation results produced by EasyText (Fig. 9). Fig. 10 and Fig. 11 further illustrate the model’s ability to generate text in multiple languages, all achieved using a single, general-purpose model. Additional figures also highlight other powerful capabilities of EasyText (Fig. 12), including layout-free text generation, adaptation to curved or slanted regions, and strong generalization to unseen characters and languages—both of which were never encountered during training.
Figure 9: Additional text rendering results generated by EasyText
Figure 10: Results of EasyText across multiple languages, showing the first group of languages.
Figure 11: Results of EasyText across multiple languages, showing the second group of languages.
Figure 12: Additional results generated by EasyText, showcasing unique capabilities beyond existing methods, such as handling unseen characters and languages.
Appendix C Additional Evaluation Details
We selected several metrics to assess the quality of text rendering results, including Image Aesthetics, Text Aesthetics, Text Quality, and Text-Image Fusion. These metrics were evaluated through two distinct evaluation methods: GPT-based automatic evaluation and a user study.
C.1 GPT-Based Evaluation
Figure 13: Example dialogue illustrating the evaluation process using GPT-4o.
The evaluation procedure for an individual image using GPT is illustrated in Fig. 13. The figure shows an example of the dialogue-based evaluation process for clarity; in practice, the evaluations were conducted by directly querying the GPT-4o API, which automatically returned the scores. Evaluations were performed at scale, and the final scores for each metric were computed as the average across all samples.
C.2 User Study Design
Table 6: User study results (%): Comparison of different methods based on user preferences across four evaluation metrics. The percentages represent the average proportion of times each method was selected as the best-performing one across multiple groups of samples.
Metrics EasyText (pretrain) EasyText (fine-tune) Glyph-v2 SD3.5 FLUX AnyText
Image Aesthetics 15.05 25.01 16.67 21.31 7.58 14.39
Font Aesthetics 15.91 31.06 9.85 15.15 13.64 14.39
Text Quality 11.65 22.97 14.39 21.10 13.96 15.91
Text-Image Integration 15.15 34.09 14.62 11.36 12.64 12.14
The user study followed a questionnaire-based format, as detailed in Fig. 14, which illustrates an example group of comparison images presented in the questionnaire. Approximately 50 questionnaires were distributed. Each questionnaire compared outputs from different models across ten groups of generated results.The results of the user study are shown in Table 6.
Figure 14: An example group of comparison images from the user study questionnaire.

### v2

@@
 Generating accurate multilingual text with diffusion models has long been desired but remains challenging.
-Recent methods have made progress in rendering text in a single language, but rendering arbitrary languages is still an unexplored area.
+Recent methods have made progress in rendering text in a single language, but rendering arbitrary languages is still an under-explored area.
 This paper introduces EasyText, a text rendering framework based on DiT (Diffusion Transformer), which connects denoising latents with multilingual character tokens encoded as character tokens.
@@
 Extensive experiments and evaluations demonstrate the effectiveness and advancement of our approach in multilingual text rendering, visual quality, and layout-aware text integration.
-Project page: https://github.com/songyiren725/EasyText
-Figure 1: Text-rendered results generated by EasyText, which supports text rendering in over ten languages and produces high-quality results.
-It can render text either with explicit positional control or in a layout-free manner, and effectively handles curved and slanted regions.
-The displayed texts do not appear in the prompts; they are included solely to illustrate the intended rendering targets.
-1 Introduction
-Scene text rendering plays a crucial role in various real-world applications.
-However, most existing methods such as TextDiffuser [7, 8], Diff-font [14] and modern commercial models like FlUX-dev [20] and Ideogram [1], are primarily limited to English, making multilingual text rendering still a challenging task.
-GlyphBy-T5-V2 [25] was the first to enable multilingual text rendering by introducing a specially designed glyph encoder and employing a multi-stage training strategy, achieving text rendering in 10 languages on a large-scale dataset of millions of images.
-Inspired by how humans learn to write, we gain several key insights: (1) Imitative writing is significantly easier than recalling—humans typically begin by mimicking (i.e., writing with visual reference) before progressing to memory-based writing.
-(2) Once familiar with one language, humans naturally develop the ability to reproduce text in other unfamiliar languages even without understanding them—treating it more like drawing than writing.
-Motivated by this, we argue that training AI to “imitate” rather than “recall” is a more efficient and effective strategy for text rendering.
-The task of scene text rendering faces several key challenges: (1) Multilingual character modeling is highly complex—for example, Chinese alone has over 30,000 characters.
-Extending to multilingual settings drastically expands the character space, making joint modeling harder, especially for rare or low-frequency characters with poor rendering accuracy.
-This is further complicated by language imbalance and font variability.
-(2) Text-background integration is often unnatural; existing methods struggle to blend rendered text with scene content, leading to visual artifacts such as disconnection or pasted-on effects that undermine image realism.
-(3) Preserving generative priors is difficult, as fine-tuning on large-scale text-image datasets improves rendering capabilities but often degrades the model’s general image generation ability.
-Table 1: Functionality evaluation of EasyText in comparison to other competitors
-Functionality Position Control Irregular Regions Multi-Lingual Long Text Rendering Text-Image Blending Unseen Characters Generalization
+Functionality Position Control Irregular Regions Multi-Lingual Long Text Rendering Text-Image Blending Unseen Characters
 Glyph-SDXL-v2
@@
 EasyText
+Table 1: Functionality evaluation of EasyText in comparison to other competitors.
+Codehttps://github.com/songyiren725/EasyText
+Datasetshttps://huggingface.co/datasets/lllrrnn/EasyText
+Extended versionhttps://arxiv.org/abs/2505.24417
+1 Introduction
+Scene text rendering is crucial for various real-world applications.
+However, most existing methods such as TextDiffuser (Chen et al.
+2023a, 2024), Diff-font (He et al.
+2024) and modern commercial models like FlUX-dev (Labs 2024) and Ideogram (Ideogram Inc.
+2025), are primarily limited to English, making multilingual text rendering still a challenging task.
+Glyph-ByT5-V2 (Liu et al.
+2024b) was one of the earliest and most representative works to enable multilingual text rendering by introducing a specially designed glyph encoder and employing a multi-stage training strategy.
+Inspired by how humans learn to write, we derive several key insights: (1) Imitative writing is considerably easier than recalling—humans typically begin by mimicking before advancing to memory-based writing.
+(2) Once familiar with one language, humans naturally develop the ability to reproduce text in other unfamiliar languages even without understanding them—treating it more like drawing than writing.
+Motivated by this, we argue that training AI to “imitate” rather than “recall” is a more efficient and effective strategy for text rendering.
+The task of scene text rendering faces several key challenges: (1) Multilingual character modeling is highly complex—for example, Chinese alone has over 30,000 characters.
+Extending to multilingual settings drastically expands the character space, making joint modeling harder, especially for rare or low-frequency characters.
+This is further complicated by language imbalance and font variability.
+(2) Text-background integration is often unnatural; existing methods struggle to blend rendered text with scene content, leading to visual artifacts such as disconnection or pasted-on effects that undermine image realism.
+(3) Preserving generative priors is difficult, as fine-tuning on large-scale text-image datasets improves rendering capabilities but often degrades the model’s general image generation ability.
 To this end, we present EasyText (shown in Fig.
-1), a novel multilingual text image generation framework based on Diffusion Transformers [30, 46].
+1), a multilingual text image generation framework based on Diffusion Transformers (Peebles and Xie 2023; Vaswani et al.
+2017).
 We encode text into font tokens via a VAE and concatenate them in the latent space with denoised latents.
-By leveraging the powerful in-context capabilities of DiT, EasyText achieves high-quality and accurate text rendering.
+Leveraging the in-context capabilities of DiT, EasyText achieves high-quality and accurate text rendering.
 Additionally, we propose a simple yet effective position control strategy called Implicit Character Position Alignment, which allows for precise control of character positions through positional encoding interpolation and replacement—enabling both position-aware rendering and layout-free generation.
-EasyText is also highly efficient in terms of data usage.
-Unlike GlyphBy-T5 V1/V2 [24, 25], which rely on contrastive synthetic data or massive real-world scene text datasets, our method simply overlays text randomly on natural images during the pretraining stage to learn glyph features.
-In this stage, to enable the model to learn glyph imitating rather than simple shape copying, we employ a multi-font mapping approach.
+EasyText is also highly data-efficient.
+Unlike Glyph-ByT5-V1/V2 (Liu et al.
+2024a, b), which rely on contrastive synthetic data or massive real-world text datasets, our method simply overlays text randomly on natural images during the pretraining stage to learn glyph features.
+To encourage the model to learn glyph imitation rather than simple shape copying, we employ a multi-font mapping approach.
 Multiple different fonts are overlaid in the synthetic training images, while the condition image uses only a standard font.
-Afterwards, we fine-tune a lightweight LoRA [16] on only 20K high-quality multilingual scene text images, enhancing the visual consistency between text and background.
+Afterwards, we fine-tune a lightweight LoRA (Hu et al.
+2022) on only 20K high-quality multilingual scene text images, enhancing the visual consistency between text and background.
 As shown in Table 1, we highlight some of the capabilities demonstrated by EasyText including: (a) Position control, precisely positioning text in specific locations within an image; (b) Irregular regions, enabling it to handle text rendering in irregular regions such as slanted or curved regions; (c) Multilingual text handling, supporting text generation in multiple languages such as Chinese, English, Japanese, etc; (d) Long text rendering, which can render extended text passages; (e) Text-image blending, which integrates text seamlessly into images, maintaining natural visual consistency; and (f) Unseen character generalization, allowing the model to exhibit generalization capability on unfamiliar and unseen characters.
 In summary, our contributions are:
-1.We propose EasyText, a framework that teaches AI to "imitate" rather than "recall," achieving high-quality multilingual text rendering by harnessing the in-context learning power of Diffusion Transformers.
-2.We introduce Implicit Character Position Alignment, which precisely controls text placement via lightweight position encoding operations and supports layout-free generation.
-3.Extensive experiments demonstrate the effectiveness and simplicity of our method, showing superior performance on challenging scenarios such as long text, multi-text layouts, irregular regions, and unseen characters.
+1.
+We propose EasyText, a framework that teaches AI to “imitate” rather than “recall”, achieving high-quality multilingual text rendering by harnessing the in-context learning power of Diffusion Transformers.
+2. We introduce Implicit Character Position Alignment, which precisely controls text placement via position encoding operations and supports layout-free generation.
+3.
+Extensive experiments demonstrate the effectiveness and simplicity of our method, showing superior performance on challenging scenarios such as long text, multi-text layouts, irregular regions, and unseen characters.
+Figure 1: Text-rendered results generated by EasyText, which supports text rendering in over ten languages and produces high-quality results.
+It can render text either with explicit positional control or in a layout-free manner, and effectively handles curved and slanted regions.
+The additional texts shown below each prompt denote the target texts to be rendered in the image which are not part of the input prompt.
 2 Related Works
 2.1 Diffusion Models
-In recent years, diffusion models [23, 35] have achieved remarkable progress and emerged as a widely used approach for image generation, offering strong capabilities in producing high-quality and diverse visual content through iterative denoising.
-Their strong capacity for modeling complex data distributions has enabled broad applicability, ranging from image synthesis and editing [4, 15] to video generation [3].
-Notable examples include Stable Diffusion [33], its enhanced version SDXL [31], and subsequent variants, which have demonstrated the scalability and effectiveness of text-to-image diffusion models for high-quality image synthesis.
-Recently, the use of transformer-based denoisers—particularly the Diffusion Transformer (DiT) [30], which replaces U-Net with a transformer backbone—has gained significant traction, and has been adopted in many state-of-the-art models such as FlUX-dev [20], Stable Diffusion 3.5 [11], and PixArt [9].
+In recent years, diffusion models (Liang et al.
+2024; Song et al.
+2020) have achieved remarkable progress and emerged as a widely used approach for image generation, offering strong capabilities in producing high-quality and diverse visual content through iterative denoising.
+Their strong capacity for modeling complex data distributions has enabled broad applicability, ranging from image synthesis and editing (Brooks et al.
+2023; Hertz et al.
+2022) to video generation (Bar-Tal et al.
+2024).
+Notable examples include Stable Diffusion (Rombach et al.
+2022), its enhanced version SDXL (Podell et al.
+2023), and subsequent variants, which have demonstrated the scalability and effectiveness of text-to-image diffusion models for high-quality image synthesis.
+Recently, the use of transformer-based denoisers—particularly the Diffusion Transformer (DiT) (Peebles and Xie 2023; Feng and Zhang 2023; Feng et al.
+2025; Feng and Zhang 2024), which replaces U-Net with a transformer backbone—has gained significant traction, and has been adopted in many state-of-the-art models such as FlUX-dev (Labs 2024), Stable Diffusion 3.5 (Esser et al.
+2024), and PixArt (Chen et al.
+2023b).
 2.2 Condition-guided Diffusion Models
-Condition-guided diffusion models aim to incorporate external signals—such as spatial structure or semantic reference—into the generative process, enabling control over layout [22, 58], content [55, 56, 50, 57, 59, 37], and identity [49].
+Condition-guided diffusion models incorporate external signals—such as spatial structure or semantic reference—into the generative process, enabling control over layout (Li et al.
+2024b; Zhang et al.
+2025), content (Zhang et al.
+2024b, c; Wang et al.
+2024b; Zhang et al.
+2024d, e; Song et al.
+2024), motion  (Ma et al.
+2025e, 2023, 2024, d, b, c), and identity (Wang et al.
+2024a).
 These methods facilitate consistent synthesis, alignment with user intent, and support for flexible customization.
 Early implementations based on U-Net architectures adopted two main paradigms: attention-based conditioning, which integrates semantic features via auxiliary encoders, and residual-based fusion, which injects spatial features into intermediate layers.
-ControlNet [54] and T2I-Adapter [28] are representative examples that improve spatial consistency and layout control under this framework.
-With the adoption of transformer-based diffusion models, recent approaches reformulate both semantic and spatial conditions as token sequences.
-These are integrated into the generation process through multi-modal attention or token concatenation mechanisms, as demonstrated in DiT-based systems like OminiControl [41, tan2025ominicontrol2] .
-This unified formulation advances the development and application of conditional image generation  [47, 38, 36, 17, 13], improving scalability, simplifying model design, and supporting efficient handling of multiple or high-resolution conditions.
+ControlNet (Zhang et al.
+2023) and T2I-Adapter (Mou et al.
+2024) are representative examples that improve spatial consistency and layout control under this framework.
+With the adoption of transformer-based diffusion models, recent approaches reformulate both semantic and spatial conditions as token sequences and integrate them into the generation process through multi-modal attention or token concatenation mechanisms, as demonstrated in DiT-based systems like OminiControl (Tan et al.
+2024) .
+This unified formulation advances the development and application of conditional image generation  (Wan et al.
+2024; Song et al.
+2025b, a; Huang et al.
+2025; Guo et al.
+2025), improving scalability, simplifying model design, and enabling efficient handling of multiple conditions.
 2.3 Visual Text Generation
-Text generation and rendering is a classic task, where early methods mainly relied on generative adversarial networks (GANs) [51, 6, 2, 42] and vector stroke techniques  [43, 39, 40].
+Text generation and rendering is a classic task, where early methods mainly relied on generative adversarial networks (GANs) (Wu et al.
+2019; Cha et al.
+2020; Azadi et al.
+2018; Tang et al.
+2022) and vector stroke techniques  (Thamizharasan et al.
+2024; Song et al.
+2023; Song and Zhang 2022).
 Recent work on visual text generation with diffusion models primarily focuses on optimizing text encoding and spatial control mechanisms to improve the fidelity and controllability of rendered text.
-Character-aware encoders are increasingly adopted: GlyphDraw [27], GlyphControl [52], and AnyText [45] embed glyph or OCR features into conditional inputs, while TextDiffuser-2 [8] leverages character-level tokenization to improve alignment.
+Character-aware encoders are increasingly adopted: GlyphDraw (Ma et al.
+2023), GlyphControl (Yang et al.
+2023), and AnyText (Tuo et al.
+2023) embed glyph or OCR features into conditional inputs, while TextDiffuser-2 (Chen et al.
+2024) leverages character-level tokenization to improve alignment.
 However, semantic bias and tokenizer limitations still persist.
 Spatial control has been addressed by introducing explicit layout-related conditions.
-TextDiffuser [7] and ControlText [18] use segmentation or layout masks, and UDiffText [60] and Brush Your Text [53] refine attention to enforce region-level alignment.
-For multilingual rendering, works like AnyText [45], GlyphControl [52], and Glyph-ByT5-v2 [25] bypass tokenizers using glyph encoders or tokenizer-free models.
+TextDiffuser (Chen et al.
+2023a) and ControlText (Jiang et al.
+2025) use segmentation or layout masks, and UDiffText (Zhao and Lian 2023) and Brush Your Text (Zhang et al.
+2024a) refine attention to enforce region-level alignment.
+For multilingual rendering, works like AnyText (Tuo et al.
+2023), GlyphControl (Yang et al.
+2023), and Glyph-ByT5-v2 (Liu et al.
+2024b) bypass tokenizers using glyph encoders or tokenizer-free models.
 Yet, rendering quality remains limited by encoder–diffusion compatibility.
-Newer models (GlyphDraw2 [26], JoyType [21], AnyText2 [44], FonTS [34], RepText [48]) improve layout precision, while closed models (Kolors 2.0 [19], Seedream 3.0 [12], GPT-4o [29]) show promising results but often suffer from limited spatial controllability and suboptimal performance in multi-text layouts.
+Newer models (GlyphDraw2 (Ma et al.
+2025a), JoyType (Li et al.
+2024a), AnyText2 (Tuo et al.
+2024), FonTS (Shi et al.
+2024), RepText (Wang et al.
+2025)) improve layout precision, while closed models (Kolors 2.0 (Kolors Team 2024), WordCon  (Shi et al.
+2025), Seedream 3.0 (Gao et al.
+2025), GPT-4o (OpenAI 2024)) show promising results but often suffer from limited spatial controllability and suboptimal performance in multi-text layouts.
 Despite progress, challenges remain in complex layouts and multilingual text, motivating further integration of glyph, position, and semantic signals.
+3 Method
 Figure 2: Overview of EasyText.
-We adopt a two-stage training strategy: large-scale pretraining for glyph generation and spatial mapping, followed by fine-tuning for visual-text integration and aesthetic refinement.
-Character positions from the condition input are aligned with target regions via implicit character position alignment, and training proceeds with image-conditioned LoRA.
-3 Method
-Section 3.1 introduces FLUX as the base model; Section 3.2 outlines the overall architecture; Section 3.3 describes the target text condition representation; Section 3.4 details the Implicit Character Position Alignment; and Section 3.5 explains the paired dataset construction.
-3.1 Preliminary: FLUX
-FLUX is an advanced generative framework significantly enhancing generation quality.
-Departing from traditional U-Net-based diffusion models (such as Stable Diffusion), it adopts a Transformer backbone (DiT) to improve long-sequence modeling and complex semantic understanding.
-The architecture employs a pre-trained T5 [32] encoder to project textual inputs into dense semantic embeddings $\mathbf{C}_{T}\in\mathbb{R}^{M\times d}$.
-For visual processing, a variational autoencoder (VAE) performs nonlinear dimensionality reduction on pixel space, generating compact latent representations that preserve essential visual semantics.
-Additionally, the framework processes encoded noise patches $\mathbf{X}\in\mathbb{R}^{N\times d}$ through rotary position embeddings (RoPE), maintaining spatial coherence during the iterative denoising process.
-These noise and text tokens are then fused by concatenation to form an image-text token sequence and then fed in parallel into a transformer-based denoising model.
-The model iteratively denoises and progressively restores a high-quality image aligned with the textual prompt.
+A two-stage training strategy is used: large-scale pretraining for glyph generation and spatial mapping, followed by fine-tuning for visual-text integration and aesthetic refinement.
+Character positions from the condition input are implicitly aligned with target regions, and training proceeds with image-conditioned LoRA.
+The training of our method is based on the open-source base model FLUX (refer to Appendix Section A for details).
+Section 3.1 outlines the overall architecture; Section 3.2 describes the target text condition representation; Section 3.3 details the Implicit Character Position Alignment; and Section 3.4 explains the paired dataset construction.
 3.1 Overall Architecture
@@
 In the second stage, we fine-tune the model on a high-quality annotated dataset to improve both the aesthetic quality of rendered text and its integration with complex scene content.
-For conditioning, the input text image is encoded into the same latent space as the target image via a VAE.
+For conditioning, the input condition image ($I_{c}$) is encoded into the same latent space as the target image ($I_{tar}$) via a VAE.
 Its positional encoding is aligned with the target region through Implicit Character Position Alignment, then concatenated with the target image’s position encoding before being passed to the FLUX DiT block.
@@
 This design philosophy stems from two critical observations:
-First, we establish a unified multilingual representation to addresses the intrinsic typographic differences between writing systems.
+First, we establish a unified multilingual representation to address the intrinsic typographic differences between writing systems.
 For alphabetic scripts (e.g., English), we use 64-pixel-high images with widths adaptive to text length, naturally capturing the connected structure of alphabetic word formations.
 This horizontal layout preserves crucial inter-character relationships and spacing conventions unique to Western typography.
-Second, for logographic systems (Chinese, Japanese, etc.), we allocate individual 64×64 square images per character.
-This configuration respects the isolated nature of ideographic characters while maintaining consistent resolution across all glyphs.
-The square aspect ratio optimally captures the balanced structure inherent to CJK characters, where stroke complexity is uniformly distributed within a square design space.
-The image-based character representations are encoded via VAE encoder into latent features.
-Subsequently, the latent representations undergo patchification through a linear projection layer, which reshapes and projects the spatial features into a sequence of conditioning tokens.
+Second, for logographic systems (Chinese, Japanese, etc.), we assign a fixed-size square image to each character.
+For Chinese characters in particular, the image size is set to 64×64 pixels.
+This configuration respects the isolated nature of ideographic characters while maintaining consistent resolution.
+The square aspect ratio optimally captures the balanced structure inherent to characters, where stroke complexity is uniformly distributed within a square design space.
+Our condition image efficiently captures the rich visual features of glyphs while providing a significantly more compact representation than many existing methods that use layout inputs matching the target image resolution.
+By including only the text to be rendered, it typically requires less than one-tenth of the spatial size, substantially reducing computational overhead.
 3.3 Implicit Character Position Alignment
-To achieve precise spatial control in font generation, we propose an Implicit Character Position Alignment (ICPA) mechanism through positional encoding extrapolation.
-Given a conditional image patch $\mathbf{P}_{c}\in\mathbb{R}^{64\times W_{c}\times 3}$ containing source typography features, $W_{c}$ denotes the total width of the conditional image patch, and a target rendering region $\Omega_{t}=[x_{1}:x_{2},y_{1}:y_{2}]$ in the output image, our method establishes position-aware feature correspondence through positional encoding extrapolation.
+To enable flexible and precise spatial control over the rendered text within the condition image, we introduce an Implicit Character Position Alignment (ICPA) mechanism, which maps the spatial coordinates of the rendered text in the target image (even in irregular regions) onto the corresponding characters in the condition image.
+Given a conditional image patch $\mathbf{P}_{c}\in\mathbb{R}^{64\times W_{c}\times 3}$ containing source typography features—where $W_{c}$ denotes the total width of the conditional image patch, and a target rendering region $\Omega_{t}$ in the output image, our method establishes position-aware feature correspondence through positional encoding extrapolation.
 Linear Alignment via Affine Transform.
-Linearly interpolate the position in the condition image (of size $64\times W_{c}$) to the position in the target bounding box $\Omega_{t}=[x_{1}:x_{2},y_{1}:y_{2}]$.
+Linearly interpolate the position in the condition image to the position in the target bounding box (an axis-aligned rectangular region in this case) $\Omega_{t}=[x_{1}:x_{2},y_{1}:y_{2}]$.
 Assuming $u\in[0,W_{c}-1]$ and $v\in[0,63]$ (i.e.
-the condition image spans indices 00 to $W_{c}$ horizontally and 00 to $63$ vertically), the affine mapping $\mathcal{T}_{\text{aff}}:(u,v)\mapsto(x,y)$ is defined as:
-$\mathcal{T}_{\text{aff}}(u,v)=\Big{(}x_{1}+\frac{u}{\,W_{c}-1\,}(x_{2}-x_{1}),\quad y_{1}+\frac{v}{\,63\,}(y_{2}-y_{1})\Big{)}\,,$ (1)
-which maps the normalized horizontal coordinate $u/(W_{c}-1)$ to a point between $x_{1}$ and $x_{2}$, and similarly $v/63$ to a point between $y_{1}$ and $y_{2}$, achieving a linear scaling and translation of the patch coordinates into the target domain.
-The positional encoding is then re-aligned by applying the affine transform to each coordinate in condition image $(u,v)$.
-Specifically, the transformed positional encoding at the target location $\text{PE}_{t}(x,y)=\mathcal{T}_{\text{aff}}(u,v)$ is assigned from the location in condition image.
-Thus, the new positional encoding for points in the condition image is given by:
-$\text{PE}_{c}(u,v)=\text{PE}_{t}(x,y),$ (2)
-where the target position $(x,y)$ is computed via the affine mapping from the patch coordinates.
+the condition image spans indices 0 to $W_{c}$ horizontally and 0 to $63$ vertically), the affine mapping $\mathcal{T}_{\text{aff}}:(u,v)\mapsto(x,y)$ is defined as:
+$\mathcal{T}_{\text{aff}}(u,v)$ $=\Big(x_{1}+\frac{u}{W_{c}-1}(x_{2}-x_{1}),$
+$\quad y_{1}+\frac{v}{63}(y_{2}-y_{1})\Big)\,$, (1)
+which maps the normalized horizontal coordinate $u/(W_{c}-1)$ to a point between $x_{1}$ and $x_{2}$, and similarly $v/63$ to a point between $y_{1}$ and $y_{2}$, achieving a linear scaling and translation of the coordinates in condition image into the target domain.
+The positional encoding in condition image is then re-aligned by applying the affine transform, yielding a transformed position $(u^{\prime},v^{\prime})=(x,y)=\mathcal{T}_{\text{aff}}(u,v)$.
+These transformed coordinates are used to update the positional encoding in condition image, enabling spatial consistency between the condition and target images.
 Nonlinear Alignment via Thin-Plate Spline Interpolation.
-For nonlinear spatial alignment, we use a Thin-Plate Spline (TPS) interpolation.
-TPS warping defines a smooth map that exactly interpolates a set of control point correspondences between the source patch and target region, while minimizing bending (second-derivative) distortion.
-Let $\{(u_{i},v_{i})\}_{i=1}^{K}$ be $K$ landmark points in the source patch $P_{c}$ and $\{(x_{i},y_{i})\}_{i=1}^{K}$ their corresponding positions in the target region $\Omega_{t}$.
+For nonlinear spatial alignment, we adopt Thin-Plate Spline (TPS) interpolation, which defines a smooth mapping that exactly fits a set of control point correspondences between the source patch and the target region, while minimizing second-order bending distortion.
+Here, $\Omega_{t}$ denotes an irregular region.
+We obtain $K$ control points along its boundary to form a deformation region that closely fits and fully covers the rendered text.
+These points act as spatial anchors for the TPS transformation.
+Let $\{(u_{i},v_{i})\}_{i=1}^{K}$ be $K$ landmark points in the condition image $P_{c}$ and $\{(x_{i},y_{i})\}_{i=1}^{K}$ their corresponding positions in the target region $\Omega_{t}$.
 The TPS mapping $\mathcal{T}_{\text{TPS}}:(u,v)\mapsto(x,y)$ can be expressed as an affine base plus a radial basis deformation term:
-$\mathcal{T}_{\text{TPS}}(u,v)=A\begin{pmatrix}u\\ v\\ 1\end{pmatrix}+\sum_{i=1}^{K}w_{i}\,\phi\!\left(\sqrt{(u-u_{i})^{2}+(v-v_{i})^{2}}\right),\quad\phi(r)=r^{2}\log r,$ (3)
+$\mathcal{T}_{\text{TPS}}(u,v)$ $=A\begin{pmatrix}u\\ v\\ 1\end{pmatrix}+\sum_{i=1}^{K}w_{i}\,\phi\!\left(\sqrt{(u-u_{i})^{2}+(v-v_{i})^{2}}\right),$
+$\phi(r)$ $=r^{2}\log r,$ (2)
 where $A\in\mathbb{R}^{2\times 3}$ represents the affine part of the transformation and $w_{i}\in\mathbb{R}^{2}$ are the TPS warp coefficients associated with each control point.
@@
 The parameters $A$ and $w_{i}$} are determined by solving the interpolation constraints
-$T_{\text{TPS}}(u_{i},v_{i})=(x_{i},y_{i}),i=1,\dots,K,$ (4)
+$\mathcal{T}_{\text{TPS}}(u_{i},v_{i})=(x_{i},y_{i}),i=1,\dots,K,$ (3)
 together with additional conditions to ensure a well-posed solution ($\sum_{i=1}^{K}w_{i}=0$, $\sum_{i=1}^{K}w_{i}u_{i}=0$, $\sum_{i=1}^{K}w_{i}v_{i}=0$).
-This yields a smooth, non-linear warp $\mathcal{T}_{\text{TPS}}$ that maps every source coordinate $(u,v)$ to a target coordinate $(x,y)$, honoring the given correspondences.
-Similar to linear alignment, target location is obtained via $\text{PE}_{t}(x,y)=\mathcal{T}_{\text{TPS}}(u,v)$.
-Accordingly, the corresponding positional encoding $\text{PE}_{c}(u,v)$ can be computed as in Equation 2.
+Similar to linear alignment, transformed position $(u^{\prime},v^{\prime})=(x,y)=\mathcal{T}_{\text{TPS}}(u,v)$ are used to update the positional encoding in condition image.
 Layout-Free Position Alignment via Positional Offset Injection.
 To enable flexible layout-free rendering, we also introduce an effective positional offset strategy.
-Instead of aligning coordinates, we shift the positional encoding of the conditional patch by a fixed scalar offset $w_{t}$, which represents the width of the target image.
-Let $(i,j)$ index the spatial position in $\mathbf{P}_{c}$.
-Target coordinates $(x^{\prime},y^{\prime})$ are obtained by applying a positional offset to the original coordinates:
-$\begin{cases}x^{\prime}=i+w_{t}\\ y^{\prime}=j\end{cases}$ (5)
-This ensures positional encoding uniqueness without binding the conditional patch to any specific target location, enabling more flexible text rendering.
+We shift the positional encoding of the conditional image by a fixed scalar offset $w_{t}$, which represents the width of the target image.
+Let $(u,v)$ index the spatial position in $\mathbf{P}_{c}$.
+Updated coordinates $(u^{\prime},v^{\prime})$ are obtained by applying a positional offset to the original coordinates:
+$\begin{cases}u^{\prime}=u+w_{t}\\ v^{\prime}=v\end{cases}.$ (4)
+This ensures positional encoding uniqueness without binding the conditional image to any specific target location, enabling more flexible text rendering.
 3.4 EasyText Dataset Construction
+Script Type Large-Scale Pre-Training Fine-Tuning
+Unique Chars Font Types Samples Samples
+Chinese 7000 18 230K 5.5K
+Korean 4308 13 120K 0.1K
+Japanese 2922 17 120K 0.1K
+Thai 2128 3 120K
+Vietnamese 91 19 120K
+Latin 104 30 180K 15K
+Greek 79 19 120K
+Table 2: Statistics of multilingual data used in large-scale pre-training and fine-tuning.
 We construct two datasets tailored to the specific needs of the pretraining and fine-tuning stages.
@@
 High-Quality Human-Annotated Dataset.
-The second dataset comprises 20k high-quality image–text box pairs, primarily in Chinese and English.
-Text regions are extracted using PP-OCR [10] and filtered to ensure annotation quality.
+The second dataset consists of approximately 20k high-quality image–text box pairs, primarily in Chinese and English.
+Text regions are extracted using PP-OCR (Du et al.
+2020) and filtered to ensure annotation quality.
 In these examples, text is more naturally integrated into scene content, often with stylized typography and visually consistent stroke patterns.
@@
 4 Experiments
+Figure 3: Qualitative comparison of EasyText with other methods, focusing on the generation quality of both text and images, reveals that EasyText demonstrates outstanding performance.
+Figure 4: Qualitative comparison of EasyText across multiple languages of the same prompt.
 4.1 Experimental Settings
@@
 Evaluation Metric.
-To evaluate the accuracy of multilingual text generation, we adopt two complementary metrics: (i) Character-level precision which measures the correctness of generated characters relative to ground-truth texts.
+To evaluate the accuracy of multilingual text generation, we adopt two complementary metrics: (i) Character-level precision, which measures the correctness of generated characters relative to ground-truth texts.
 This reflects the model’s ability to produce accurate predictions at the most basic unit level for each language.
@@
 Multilingual Benchmark.
-We construct a multilingual benchmark with 90 language-agnostic prompts across 10 languages.
+We construct a multilingual benchmark with 90 language-agnostic prompts.
 For each language, prompts are paired with language-specific text to generate condition images while preserving semantic intent.
 Four images are generated per prompt-language pair.
-For logographic scripts (Chinese, Japanese, Korean), the average number of characters per text box is 7.1, and for alphabetic scripts, it is 14.3.
+For logographic scripts, the average number of characters per text box is 7.1, and for alphabetic scripts, it is 14.3.
 Each image contains an average of 1.7 text boxes.
 To evaluate the generated results, three volunteers were tasked with assessing the accuracy of the text in the images, ensuring reliable multilingual generation quality.
-Table 2: Statistics of multilingual data used in large-scale pre-training and fine-tuning.
-Script Type Large-Scale Pre-Training Fine-Tuning
-Unique Chars Font types Samples Samples
-Chinese 7000 18 230K 5.5K
-Korean 4308 13 100K 0.1K
-Japanese 2922 17 100K 0.1K
-Thai 2128 3 100K
-Vietnamese 91 19 100K
-Latin 104 30 180K 16K
-Greek 79 19 100K
-Table 3: Character/Sentence level precision for English and Chinese across different methods.
+4.2 Comparison Results
 Method English Chinese
@@
 ANYTEXT 0.8978 0.6364 0.8824 0.6071
-Table 4: Multilingual text generation precision with our unified and generalized model, compared to Glyph-SDXL-v2.
+Table 3: Character/Sentence level precision for English and Chinese across different methods.
+We evaluate our model on a multilingual benchmark against state-of-the-art commercial models (e.g., FLUX (Labs 2024), Jimeng AI (ByteDance 2024), SD3.5 (Esser et al.
+2024)) and other methods (e.g., Glyph-ByT5-v2 (Liu et al.
+2024b), AnyText (Tuo et al.
+2023)), following the benchmark and metrics outlined in Sec.
+4.1.
+For languages with limited support in prior rendering methods, we compare primarily with Glyph-SDXL-v2.
+All baseline models are evaluated using their official inference settings.
+For each sample, four images are generated with identical image descriptions and target texts across all methods.
+For models lacking conditional input support, the target text is included in the prompt; evaluation focuses solely on text accuracy, regardless of spatial layout or surrounding content.
+Comprehensive Quality Assessment.
+Beyond precision evaluation, we assess generation quality using CLIPScore and OCR accuracy for objective comparison against existing baselines.
+We further incorporate GPT-4o assessment and a user study across four subjective criteria: Image Aesthetics, Text Aesthetics, Text Quality, and Text-Image Fusing, to evaluate overall fidelity and alignment.
+The results show that our model surpasses competing methods in text rendering precision across several languages, including English and Italian (Table 4) and achieves strong character-level accuracy in Chinese (Table 3), though slightly behind Jimeng AI at the sentence level.
+It also performs well in unsupported languages like Thai and Greek.
+In Table 5, it also leads in OCR accuracy and improves over FLUX in CLIPScore, indicating higher visual-text alignment.
+These trends are further supported by GPT-4o evaluation results.
+After pretraining, the model demonstrates strong text rendering performance from condition images.
+However, it shows limited text-image coherence, as indicated by lower CLIPScore and GPT-4o evaluations.
+Fine-tuning alleviates this issue, significantly improving CLIPScore, Text Aesthetics, and overall visual-text alignment.
 Language EasyText (pre) EasyText (fine) Glyph-ByT5-v2
@@
 Greek 0.9702 0.7685 0.9360 0.6875
-Table 5: Quantitative evaluation of text-to-image rendering across diverse metrics and GPT-4o-based assessments.Ours (pre) and Ours (ft) denote the pretrained and finetuned versions of our model, respectively.
-Metric Ours(pre) Ours(ft) Glyph-v2 SD3.5 FLUX
-CLIPScore 0.3318 0.3486 0.3140 0.3519 0.3348
-OCR Acc.(%) 84.32 88.72 82.33 79.33 76.09
-Evaluation by GPT-4o
-Image Aesthetics 81.58 83.86 75.86 84.58 83.61
-Text Aesthetics 65.14 73.79 65.28 72.06 71.90
-Text Quality 84.58 90.66 86.25 85.79 86.95
-Text-Image Fusing 74.48 80.28 74.80 81.12 78.66
-Figure 3: Qualitative comparison of EasyText with other methods, focusing on the generation quality of both text and images, reveals that EasyText demonstrates outstanding performance.
-Figure 4: Qualitative comparison of EasyText across multiple languages of the same prompt.
-4.2 Comparision results
-Multi-lingual Precision Evaluation.
-We evaluate our model on a multilingual benchmark against state-of-the-art commercial models (e.g., FLUX [20], Jimeng AI [5], SD3.5 [11]) and other methods (e.g., Glyph-ByT5-v2 [25], AnyText [45]), following the benchmark and metrics outlined in Sec.
-4.1.
-For languages with limited support in prior rendering methods, we compare primarily with Glyph-SDXL-v2.
-All baseline models are evaluated using their official inference settings.
-For each sample, four images are generated with identical image descriptions and target texts across all methods.
-For models lacking conditional input support, the target text is included in the prompt; Evaluation focuses solely on text accuracy, regardless of spatial layout or surrounding content.
-Figure 5: Comparison of text rendering on the pretraining synthetic dataset using multiple font overlays versus single font overlays. The red boxes highlight the erroneous regions.
-Comprehensive Quality Assessment.
-Beyond precision evaluation, we assess generation quality using CLIPScore and OCR Accuracy for objective comparison against existing baselines.
-We further incorporate GPT-4o assessment and a user study across four subjective criteria: Image Aesthetics, Text Aesthetics, Text Quality, and Text-Image Fusing, to evaluate overall fidelity and alignment.
-The results show that our model surpasses competing methods in text rendering precision across several languages, including English and Italian (Table 5) and achieves strong character-level accuracy in Chinese, though slightly behind Jimeng AI at the sentence level.
-It also performs well in unsupported languages like Thai and Greek.
-In Table 5, it also leads in OCR accuracy and improves over FLUX in CLIPScore, indicating higher visual-text alignment.
-These trends are further supported by GPT-4o evaluation results.
-After pretraining, the model demonstrates strong text rendering performance from condition images.
-However, it shows limited text-image coherence, as indicated by lower CLIPScore and GPT-4o evaluations.
-Fine-tuning alleviates this issue, significantly improving CLIPScore, Text Aesthetics, and overall visual-text alignment.
-Although rendering accuracy slightly declines due to flexible positioning and artistic font usage, the trade-off is acceptable given the qualitative gains.
-Notably, while most competing methods use datasets exceeding 1M samples, our fine-tuning relies on only  20k annotated examples.
-Moreover, the large-scale pretraining set can be replaced with synthetic data, reducing dependence on large, annotated datasets.
+Table 4: Multilingual text generation precision with our unified and generalized model, compared to Glyph-SDXL-v2.
 4.3 Qualitative Results
@@
 Specifically, our approach generalizes well to unseen characters, slanted or non-rectilinear text regions, and long text sequences, while maintaining structural consistency and legibility.
-Detailed qualitative results are provided in the appendix.
-Figure 6: Ablation study comparing the full conditioned method with: (1) layout-free (2) without condition inputs, and (3) original FLUX. The red boxes highlight the erroneous regions.
+(refer to Appendix Section C for details)
+Metric Ours (pre) Ours (ft) Glyph-v2 SD3.5 FLUX
+CLIPScore 0.3318 0.3486 0.3140 0.3519 0.3348
+OCR Acc. (%) 84.32 88.72 82.33 79.33 76.09
+Evaluation by GPT-4o
+Image Aesthetics 81.58 83.86 75.86 84.58 83.61
+Text Aesthetics 65.14 73.79 65.28 72.06 71.90
+Text Quality 84.58 90.66 86.25 85.79 86.95
+Text-Image Fusing 74.48 81.28 74.80 81.12 78.66
+Table 5: Quantitative evaluation of text-to-image rendering across diverse metrics and GPT-4o-based assessments.
+Ours (pre) and Ours (ft) denote the pretrained and finetuned versions of our model, respectively.
 4.4 Ablation Studies
+Figure 5: Ablation study comparing the full conditioned method with: (1) layout-free (2) without condition inputs, and (3) original FLUX-Dev.
+<sks1>, <sks2> triggers are used to replace rendered text in the first two cases.
 To validate the effectiveness of using diverse fonts in synthetic pretraining, we conducted a controlled experiment where both the target and condition images were rendered using the same standard font.
-While this setup yields high initial precision, it mainly captures a direct shape mapping and fails to generalize after fine-tuning on real-world data(Shown in Fig.
-5).
-This highlights the importance of font diversity—rendering targets with multiple fonts while keeping the condition font fixed—for learning robust, transferable representations.
-we compare our full approach—using both the condition input and implicit character position alignment-with three ablated variants: (1) removing position alignment, (2) removing the condition input, and (3) using the original FLUX model.
+Despite high initial precision, this setup captures only direct shape mappings and fails to generalize after real-world fine-tuning.
+(refer to Appendix Section E) This highlights the importance of font diversity—rendering targets with multiple fonts while keeping the condition font fixed—for learning robust, transferable representations.
+We compare our approach with three ablated variants: (1) removing position alignment, (2) removing the condition input, and (3) using the original FLUX model.
 As shown in Fig.
-6, incorporating position mapping substantially improves spatial precision and text accuracy, particularly in multi-text scenarios.
-In comparison, when removing the condition input, the target render text is instead provided in the prompt(in our full method, renter text content is given via the condition input).
+5, incorporating position mapping substantially improves spatial precision and text accuracy, particularly in multi-text scenarios.
+In comparison, when removing the condition input, the target rendered text is instead provided in the prompt(in our full method, rendered text content is given via the condition input).
 This demonstrates that our fine-tuned model effectively retains the strong generation capability of the base FLUX model.
@@
 The Implicit Character Position Alignment mechanism is less effective when character positions are substantially overlapping, occasionally leading to reduced rendering accuracy.
-In addition, training across multiple scripts results in confusions between simple but visually similar characters from different writing systems, such as rendering the Vietnamese character "" as "e".
+In addition, training across multiple scripts results in confusion between simple but visually similar characters from different writing systems.
 These cases are infrequent but consistently observed.
@@
 In this work, we propose EasyText, a diffusion-based framework for multilingual and controllable text generation in text-to-image synthesis.
-The method learns to mimic font rendering by mapping glyph features into the same latent space as the image via a VAE, and achieves Implicit Character Position Alignment by combining position encoding with spatial targets, enabling precise and flexible text placement.
-Benefiting from a two-stage training strategy—pretraining on a large-scale synthetic multi-font dataset followed by fine-tuning—EasyText reduces the need for real multilingual data while maintaining high rendering accuracy.
+EasyText learns to mimic glyph features and supports precise and flexible text placement through implicit character alignment.
+With a two-stage training strategy, the method significantly reduces reliance on real multilingual data while maintaining high rendering accuracy.
 It also demonstrates strong visual-text integration, effectively embedding text into complex scenes.
-Extensive experiments confirm the effectiveness and generalization ability of our approach.
-Supplementary Material for EasyText: Controllable Diffusion Transformer for Multilingual Text Rendering
-David S. HippocampusUse footnote for providing further information about author (webpage, alternative address)—not for acknowledging funding agencies.
-Department of Computer Science
-Cranberry-Lemon University
-Pittsburgh, PA 15213
-hippo@cs.cranberry-lemon.edu
-Appendix
-This supplementary material is organized into three sections.
-Section A describes the training datasets used for pretraining and fine-tuning, including illustrative examples.
-Section 12 presents diverse and comprehensive generated examples.
-Section C outlines the process and sample cases of the GPT-based evaluation and user study.
-Appendix A Additional Dataset Details
-The main paper provides a detailed description of the dataset’s design and composition.
-This section complements that discussion by offering additional details and presenting example instances from the dataset to further illustrate its structure.
-Figure 7: Sample images from the synthetic dataset used in the pretraining.
-A.1 Synthetic Dataset for Large-Scale Pretraining
-For the training dataset used in pretraining, we collected background images from several publicly available datasets without textual content, such as LAION-Aesthetics.
-For each script (i.e., a language-specific character set), we rendered text in various fonts onto these backgrounds.
-During the rendering process, we recorded both the text content and the corresponding bounding box coordinates, producing synthetic image–text pairs designed for text rendering tasks.
-This approach enabled the creation of a large-scale multilingual dataset of approximately 1 million samples, offering a practical alternative for low-resource languages, where high-quality annotated image–text pairs are often unavailable.
-Some sample instances are shown in Fig.
-7.
-A.2 High-Quality Dataset for Fine-Tuning
-For the fine-tuning stage, we collected around 20K images containing text from copyright-free websites.
-Text regions were detected and filtered using PP-OCR, and each image was further annotated with a caption generated by GPT-4 Mini.
-Some sample instances are shown in Fig.
-8.
-Figure 8: Sample images from the high-quality dataset used in the fine-tuning.
-Appendix B Additional Generation Results
-We present additional diverse generation results produced by EasyText (Fig.
-9).
-Fig.
-10 and Fig.
-11 further illustrate the model’s ability to generate text in multiple languages, all achieved using a single, general-purpose model.
-Additional figures also highlight other powerful capabilities of EasyText (Fig.
-12), including layout-free text generation, adaptation to curved or slanted regions, and strong generalization to unseen characters and languages—both of which were never encountered during training.
-Figure 9: Additional text rendering results generated by EasyText
-Figure 10: Results of EasyText across multiple languages, showing the first group of languages.
-Figure 11: Results of EasyText across multiple languages, showing the second group of languages.
-Figure 12: Additional results generated by EasyText, showcasing unique capabilities beyond existing methods, such as handling unseen characters and languages.
-Appendix C Additional Evaluation Details
-We selected several metrics to assess the quality of text rendering results, including Image Aesthetics, Text Aesthetics, Text Quality, and Text-Image Fusion.
-These metrics were evaluated through two distinct evaluation methods: GPT-based automatic evaluation and a user study.
-C.1 GPT-Based Evaluation
-Figure 13: Example dialogue illustrating the evaluation process using GPT-4o.
-The evaluation procedure for an individual image using GPT is illustrated in Fig.
-13.
-The figure shows an example of the dialogue-based evaluation process for clarity; in practice, the evaluations were conducted by directly querying the GPT-4o API, which automatically returned the scores.
-Evaluations were performed at scale, and the final scores for each metric were computed as the average across all samples.
-C.2 User Study Design
-Table 6: User study results (%): Comparison of different methods based on user preferences across four evaluation metrics.
-The percentages represent the average proportion of times each method was selected as the best-performing one across multiple groups of samples.
-Metrics EasyText (pretrain) EasyText (fine-tune) Glyph-v2 SD3.5 FLUX AnyText
-Image Aesthetics 15.05 25.01 16.67 21.31 7.58 14.39
-Font Aesthetics 15.91 31.06 9.85 15.15 13.64 14.39
-Text Quality 11.65 22.97 14.39 21.10 13.96 15.91
-Text-Image Integration 15.15 34.09 14.62 11.36 12.64 12.14
-The user study followed a questionnaire-based format, as detailed in Fig.
-14, which illustrates an example group of comparison images presented in the questionnaire.
-Approximately 50 questionnaires were distributed.
-Each questionnaire compared outputs from different models across ten groups of generated results.The results of the user study are shown in Table 6.
-Figure 14: An example group of comparison images from the user study questionnaire.